what is the x-coordinate of the x-intercept of the function 2/3y=6x+12?
Answer by Maths68(1474)

x-intercept is a point on the graph where y=0
and
y-intercept is a point on the graph where x=0
So for the given function if we put y=0
2/3y=6x+12
2/3(0)=6x+12
0=6x+12
-6x=12
-6x/-6=12/-6
x=-2
x-intercept of the given function (-2,0)